OpenStudy (anonymous):

could someone please show me how to find f^-1 of y=x^2+6x+7 algebraicly?

OpenStudy (zarkon):

it does not have an inverse unless you restrict the domain

OpenStudy (dumbcow):

solve for x use completing the square

myininaya (myininaya):

right! \[y=(x^2+6x)+7=(x^2+6x+9)+7-9=(x+3)^2-2 \] vertex is (-3,-2) so say we restrict the domain to (-3,inf) so we have a 1 to 1 function now \[0=x^2+6x+7-y\] \[x=\frac{-6 \pm \sqrt{36-4(7-y)}}{2}=\frac{-6 \pm \sqrt{4}\sqrt{9-(7-y)}}{2}\] \[=-3 \pm \sqrt{2+y}\] so we have either \[f^{-1}(x)=-3+\sqrt{2+y}\] or \[f^{-1}(x)=-3-\sqrt{2+y}\] but not both to determine which one it is we need to think our function has domain (-3,inf) so then domain of f inverse should have range (-3,inf) so its the first once like shoose a number like 7 plug it in for the first one we get 0 the second one gives us -6 only the first is in the range so \[f^{-1}(x)=-3 + \sqrt{2+y}\]

OpenStudy (anonymous):

oooh! completing the square! that was why I couldn't solve the equation

myininaya (myininaya):

oops so then domain of f inverse should have range (-3,inf)<this sentence is weird it should say the range of f inverse is (-3,inf)

myininaya (myininaya):

shoose lol should be choose

OpenStudy (dumbcow):

wait i think its other way around domain of f: All real range of f: (2,inf) domain of f^-1 (2,inf) range of f^-1; All real

OpenStudy (dumbcow):

*(-2,inf)

OpenStudy (zarkon):

if f is defined on all of R then f^{-1} does not exist

myininaya (myininaya):

what i was doing above cow was restricting the domain of f such that the function would be 1 to 1 and therefore have an inverse

OpenStudy (dumbcow):

ok then restrict domain of f to (-3,inf) then range of f^-1 is (-3,inf) which is what myininaya was saying

myininaya (myininaya):

lol

myininaya (myininaya):

its like you knew what i was going to say

OpenStudy (dumbcow):

yeah i got mixed up a little bit